# Modulo arithmetic of function spaces: Subset hyperspaces as quotients of function spaces

### v1

@@ 5. Quotient-realization of hyperspace topologies
Definition 5.1 ($q$-lifts of subset hyperspace topologies).
@@
Remark 5.2 (Existence of $q$-lifts, Smallest $q$-lift).
@@
Remark 5.3 (Temporary notation for the proof of Theorem 5.4).
@@
    In the restriction $q:(\mathcal{I},\tau_{i})\subset X^{Y}\rightarrow(q(\mathcal{I}),\tau_{iq})\subset Cl_{Y}(X)$, if $B\subset Cl_{Y}(X)$, then we write “$B\in\tau_{iq}$” to mean “$B\cap q(\mathcal{I})\in\tau_{iq}$”.
@@
    In the restrictions $q:(\mathcal{I},\tau_{i})\rightarrow(q(\mathcal{I}),\tau_{iq})$ and $q:(\mathcal{J},\tau_{j})\rightarrow(q(\mathcal{J}),\tau_{jq})$, we write “$\tau_{iq}\subset\tau_{jq}$” to mean “$\forall B\subset Cl_{Y}(X)$, $B\in\tau_{iq}$Rightarrow $B\in\tau_{jq}$ (i.e., $B\cap q(\mathcal{I})\in\tau_{iq}$Rightarrow $B\cap q(\mathcal{J})\in\tau_{jq}$)”.
@@
Let $X$ be a space, $Y$ a set, $\mathcal{F}\subset X^{Y}$ a $q$-full finitely $q$-stable subset, and $\tau_{0}\supset\tau_{v}$ a topology on $Cl(X)$. Further suppose that some finite subset space $FS_{n}(X)\subset(Cl_{Y}(X),\tau_{v})$ is closed. Then there exists a topology $\widetilde{\tau}_{0}\supset\tau_{p}$ on $\mathcal{F}$, which is maximal with respect to subsets of $\mathcal{F}$, such that $\tau_{v}\subset\widetilde{\tau}_{0q}\subset\tau_{0}$ in $Cl_{Y}(X)$.
@@
By Theorem 3.17(1) and Remark 3.18(1), $\mathcal{P}\neq\emptyset$ (because with $1\leq n\leq|Y|$, $\mathcal{A}:=\{f\in\mathcal{F}:|q(f)|\leq n\}=q^{-1}(FS_{n}(X))\cap\mathcal{F}$, and $\tau:=\tau_{p}$, we get $\tau_{q}=\tau_{v}$ in $q:(\mathcal{A},\tau)\rightarrow(q(\mathcal{A}),\tau_{q})\subset FS_{n}(X)$). Given a chain $\{(\mathcal{A}_{\lambda},\tau_{\lambda})\}_{\lambda\in\Lambda}$ in $\mathcal{P}$, let $\mathcal{A}:=\bigcup_{\lambda}\mathcal{A}_{\lambda}$ and $\tau:=\bigcup_{\lambda}\tau_{\lambda}$ (i.e., for any $A\subset\mathcal{A}$, we have $A\in\tau$ $\iff$ $A\cap\mathcal{A}_{\lambda}\in\tau_{\lambda}$ for each $\lambda$). Then (i) $\mathcal{A}\subset\mathcal{F}$, (ii) $\tau_{p}\subset\tau$, and (iii) $\tau_{v}\subset\tau_{q}\subset\tau_{0}$ (since $\tau_{v}\subset\tau_{\lambda q}\subset\tau_{0}$ for each $\lambda$ and $\tau_{q}=\bigcap_{\lambda}\tau_{\lambda q}$). Here, $\tau_{q}=\bigcap_{\lambda}\tau_{\lambda q}$ is due to the following: Given $B\in Cl_{Y}(X)$, for each $\lambda$, $B\in\tau_{\lambda q}$ $\stackrel{{\scriptstyle\textrm{def.}}}{{\iff}}$ $q^{-1}(B)\cap\mathcal{A}_{\lambda}\in\tau_{\lambda}$, and so $B\in\tau_{q}$ $\iff$ $q^{-1}(B)\in\tau$ $\stackrel{{\scriptstyle\textrm{def.}}}{{\iff}}$ $q^{-1}(B)\cap\mathcal{A}_{\lambda}\in\tau_{\lambda}$ $\forall\lambda$ $\iff$ $B\in\tau_{\lambda q}$ $\forall\lambda$.
@@
Therefore $(\mathcal{A},\tau)$ is an upper bound of the given chain in $\mathcal{P}$. By Zorn’s lemma, $\mathcal{P}$ has a maximal element $(\mathcal{A}^{\prime},\tau^{\prime})$. Suppose $\mathcal{A}^{\prime}\neq\mathcal{F}$. Let $f\in\mathcal{F}\backslash\mathcal{A}^{\prime}$, $\mathcal{A}^{\prime\prime}:=\mathcal{A}^{\prime}\cup\{f\}$, and $\tau^{\prime\prime}$ be the topology on $\mathcal{A}^{\prime\prime}$ given by $\tau^{\prime\prime}:=\{A\subset\mathcal{A}^{\prime\prime}:A\cap\mathcal{A}^{\prime}\in\tau^{\prime}\}=\tau^{\prime}\cup\big{\{}A\cup\{f\}:A\in\tau^{\prime}\big{\}}$.
@@
We claim that $(\mathcal{A}^{\prime},\tau^{\prime})<\big{(}\mathcal{A}^{\prime\prime},\tau^{\prime\prime}\big{)}\in\mathcal{P}$. Proof of claim: It is clear that $\mathcal{A}^{\prime}\subsetneq\mathcal{A}^{\prime\prime}$ and $\tau^{\prime}\subset\tau^{\prime\prime}$ (which implies $\tau^{\prime}_{q}\supset\tau_{q}^{\prime\prime}$ as before), and so $(\mathcal{A}^{\prime},\tau^{\prime})<\big{(}\mathcal{A}^{\prime\prime},\tau^{\prime\prime}\big{)}$. Next, $\big{(}\mathcal{A}^{\prime\prime},\tau^{\prime\prime}\big{)}\in\mathcal{P}$ follows from (i) $\mathcal{A}^{\prime\prime}\subset\mathcal{F}$, (ii) $\tau_{p}\subset\tau^{\prime\prime}$, and (iii) $\tau_{v}\subset\tau_{q}^{\prime}=\tau_{q}^{\prime\prime}\subset\tau_{0}$, since we also have $\tau_{q}^{\prime}\subset\tau_{q}^{\prime\prime}$ (because given $B\in Cl_{Y}(X)$, $B\in\tau_{q}^{\prime}$ $\iff$ $q^{-1}(B)\cap\mathcal{A}^{\prime}=(q^{-1}(B)\cap\mathcal{A}^{\prime\prime})\cap\mathcal{A}^{\prime}\in\tau^{\prime}$ $\stackrel{{\scriptstyle\textrm{construction}}}{{\Longrightarrow}}$ $q^{-1}(B)\cap\mathcal{A}^{\prime\prime}\in\tau^{\prime\prime}$ $\iff$ $B\in\tau_{q}^{\prime\prime}$). This completes the proof of the claim.
@@ 6. Conclusion and questions
For a space $X$, we have seen (Theorem 3.17) that certain hyperspaces $\mathcal{J}\subset Cl(X)$ of $X$ can be described as quotients of function spaces $\mathcal{F}\subset X^{Y}$ in a natural way. Following this, we have discussed the concrete realization of certain preferred function space topologies (Theorem 4.4), the metrization of compact subset hyperspaces (Theorems 4.8,4.12), and the existence of a $\tau_{p}$-compatible function space topologies with $\tau_{v}$-compatible quotients (Theorem 5.4).
In addition to Questions 3.8,5.5, we have the following interesting questions.
Question 6.1.
Let $X,Y$ be spaces, $\mathcal{F}\subset X^{Y}$ a $q$-full subset, $Z\subset Y$, and $Cl_{Z}(X)\subset(Cl_{Y}(X),\tau)$ a $\tau$-closed subset, in which case a quotient map $q:(\mathcal{F},\widetilde{\tau})\rightarrow(Cl_{Y}(X),\tau)$ restricts to a quotient map $q:(\mathcal{F}_{Z},\widetilde{\tau})\subset X^{Y}\rightarrow(Cl_{Z}(X),\tau)$, where $\mathcal{F}_{Z}:=q^{-1}(Cl_{Z}(X))\cap\mathcal{F}(Y,X)$. Can we find a Lipschitz retraction
@@
Question 6.2.
Let $X$ be a space and $Y$ a set. From Definition 3.1, for which $Y$ (of smallest possible cardinality) does the equality $Cl_{Y}(X)=Cl(X)$ hold? Also, what is the cardinally-smallest $q$-full subset $\mathcal{F}\subset X^{Y}$?
Question 6.3 (Path representation problem).
Let $X$ be a space, $Y$ a set, $\mathcal{F}\subset X^{Y}$ a $q$-full subset, and $\tau_{\pi}$ a swrc-topology on $\mathcal{F}$. It is clear that a path $\eta:[0,1]\rightarrow(\mathcal{F},\tau_{\pi})$ gives a path $q\circ\eta:[0,1]\stackrel{{\scriptstyle\eta}}{{\longrightarrow}}\mathcal{F}\stackrel{{\scriptstyle q}}{{\longrightarrow}}(Cl_{Y}(X),\tau_{\pi q})$, since the composition of continuous maps is continuous. Conversely, when does every path $\gamma:[0,1]\rightarrow(Cl_{Y}(X),\tau_{\pi q})$ come from (or lift as $\gamma=q\circ\eta$ to) a path $\eta:[0,1]\rightarrow(\mathcal{F},\tau_{\pi})$?
Whenever the answer to this question is positive, every path $\gamma:[0,1]\rightarrow(Cl_{Y}(X),\tau_{\pi q})$ is expressible in the form $\gamma(t)=cl_{X}{\{\gamma_{r}(t):r\in\Gamma\}}$, $\forall t\in[0,1]$, for a set of paths $\{\gamma_{r}:[0,1]\rightarrow X\}_{r\in\Gamma}$ in $X$. Of course, by the axiom of choice, any path $\gamma:[0,1]\rightarrow(Cl_{Y}(X),\tau_{\pi q})$ can be written as
@@
By the above paragraph, Question 6.3 is relevant to the “path representation” problem considered in [akofor2024] and therefore relevant to [akofor2024, Question 5.1], even though the current topology $\tau_{\pi q}$ on $BCl_{Y}(X)\subset Cl_{Y}(X)$ is in general not the same as the $d_{H}$-topology. Since $\gamma^{-1}(\mathcal{O})=\eta^{-1}(q^{-1}(\mathcal{O}))$ for any open set $\mathcal{O}\subset(Cl_{Y}(X),\tau_{\pi q})$, if $\tau_{\pi}=q^{-1}(\tau_{\pi q})$, then $\gamma$ is continuous iff $\eta$ is continues (but in general, if we have proper containment $q^{-1}(\tau_{\pi q})\subsetneq\tau_{\pi}$, then $\gamma=q\circ\eta$ can be continuous even when $\eta$ is not continuous). Therefore, for $\gamma=q\circ\eta$ to be continuous, we only need a $q^{-1}(\tau_{\pi q})$-continuous (not necessarily a $\tau_{\pi}$-continuous) selection

### v2

@@ 5. Quotient-realization of hyperspace topologies
Definition 14 ($q$-lifts of subset hyperspace topologies).
@@
Remark (Existence of $q$-lifts, Smallest $q$-lift).
@@
Remark (Temporary notation for the proof of Theorem 5.1).
@@
    In the restriction $q:(\mathcal{I},\tau_{i})\subset X^{Y}\rightarrow(q(\mathcal{I}),\tau_{iq})\subset Cl_{Y}(X)$, if $B\subset Cl_{Y}(X)$, then we write $B\in\tau_{iq}$” to mean $B\cap q(\mathcal{I})\in\tau_{iq}$”.
@@
    In the restrictions $q:(\mathcal{I},\tau_{i})\rightarrow(q(\mathcal{I}),\tau_{iq})$ and $q:(\mathcal{J},\tau_{j})\rightarrow(q(\mathcal{J}),\tau_{jq})$, we write $\tau_{iq}\subset\tau_{jq}$” to mean $\forall B\subset Cl_{Y}(X)$, $B\in\tau_{iq}$ $\Rightarrow$ $B\in\tau_{jq}$ (i.e., $B\cap q(\mathcal{I})\in\tau_{iq}$ $\Rightarrow$ $B\cap q(\mathcal{J})\in\tau_{jq}$)”.
@@
Let $X$ be a space, $Y$ a set, $\mathcal{F}\subset X^{Y}$ a $q$-full finitely $q$-stable subset, and $\tau_{0}\supset\tau_{v}$ a topology on $Cl(X)$. Further suppose that some finite-subset hyperspace $FS_{n}(X)\subset(Cl_{Y}(X),\tau_{v})$ is closed. Then there exists a topology $\widetilde{\tau}_{0}\supset\tau_{p}$ on $\mathcal{F}$, which is maximal with respect to subsets of $\mathcal{F}$, such that $\tau_{0}\supset\widetilde{\tau}_{0q}\supset\tau_{v}$ in $Cl_{Y}(X)$.
@@
By Theorem 3.3(1) and Theorem 3.3’s Remark (1), $\mathcal{P}\neq\emptyset$ (because with $1\leq n\leq|Y|$, $\mathcal{A}:=\{f\in\mathcal{F}:|q(f)|\leq n\}=q^{-1}(FS_{n}(X))\cap\mathcal{F}$, and $\tau:=\tau_{p}$, we get $\tau_{q}=\tau_{v}$ in $q:(\mathcal{A},\tau)\rightarrow(q(\mathcal{A}),\tau_{q})\subset FS_{n}(X)$). Given a chain $\{(\mathcal{A}_{\lambda},\tau_{\lambda})\}_{\lambda\in\Lambda}$ in $\mathcal{P}$, let $\mathcal{A}:=\bigcup_{\lambda}\mathcal{A}_{\lambda}$ and $\tau:=\bigcup_{\lambda}\tau_{\lambda}$ (i.e., for any $A\subset\mathcal{A}$, we have $\in\tau \iff A𝒜$$\cap\mathcal{A}_{\lambda}\in\tau_{\lambda}$ for each $\lambda$). Then (i) $\mathcal{A}\subset\mathcal{F}$, (ii) $\tau_{p}\subset\tau$, and (iii) $\tau_{v}\subset\tau_{q}\subset\tau_{0}$ (since $\tau_{v}\subset\tau_{\lambda q}\subset\tau_{0}$ for each $\lambda$ and $\tau_{q}=\bigcap_{\lambda}\tau_{\lambda q}$). Here, $\tau_{q}=\bigcap_{\lambda}\tau_{\lambda q}$ is due to the following: Given $B\in Cl_{Y}(X)$, for each $\lambda$, $B\in\tau_{\lambda q}$ $\stackrel{{\scriptstyle\textrm{def.}}}{{\iff}}$ $q^{-1}(B)\cap\mathcal{A}_{\lambda}\in\tau_{\lambda}$, and so $B\in\tau_{q}$ $\iff$ $q^{-1}(B)\in\tau$ $\stackrel{{\scriptstyle\textrm{def.}}}{{\iff}}$ $q^{-1}(B)\cap\mathcal{A}_{\lambda}\in\tau_{\lambda}$ $\forall\lambda$ $\iff$ $B\in\tau_{\lambda q}$ $\forall\lambda$.
@@
Therefore $(\mathcal{A},\tau)$ is an upper bound of the given chain in $\mathcal{P}$. By Zorn’s lemma, $\mathcal{P}$ has a maximal element $(\mathcal{A}^{\prime},\tau^{\prime})$. Suppose $\mathcal{A}^{\prime}\neq\mathcal{F}$. Let $f\in\mathcal{F}\backslash\mathcal{A}^{\prime}$, $\mathcal{A}^{\prime\prime}:=\mathcal{A}^{\prime}\cup\{f\}$, and $\tau^{\prime\prime}$ be the topology on $\mathcal{A}^{\prime\prime}$ given by $\tau^{\prime\prime}:=\{A\subset\mathcal{A}^{\prime\prime}:A\cap\mathcal{A}^{\prime}\in\tau^{\prime}\}=\tau^{\prime}\cup\big\{A\cup\{f\}:A\in\tau^{\prime}\big\}$ (footnote [^13] This is the superspace topology defined in the footnote(s) on page Remark.).
@@
We claim that $(\mathcal{A}^{\prime},\tau^{\prime})<\big(\mathcal{A}^{\prime\prime},\tau^{\prime\prime}\big)\in\mathcal{P}$. Proof of claim: It is clear that $\mathcal{A}^{\prime}\subsetneq\mathcal{A}^{\prime\prime}$ and $\tau^{\prime}\subset\tau^{\prime\prime}$ (which implies $\tau^{\prime}_{q}\supset\tau_{q}^{\prime\prime}$ as before), and so $(\mathcal{A}^{\prime},\tau^{\prime})<\big(\mathcal{A}^{\prime\prime},\tau^{\prime\prime}\big)$. Next, $\big(\mathcal{A}^{\prime\prime},\tau^{\prime\prime}\big)\in\mathcal{P}$ follows from (i) $\mathcal{A}^{\prime\prime}\subset\mathcal{F}$, (ii) $\tau_{p}\subset\tau^{\prime\prime}$, and (iii) $\tau_{v}\subset\tau_{q}^{\prime}=\tau_{q}^{\prime\prime}\subset\tau_{0}$, since we also have $\tau_{q}^{\prime}\subset\tau_{q}^{\prime\prime}$ (because given $B\in Cl_{Y}(X)$, $B\in\tau_{q}^{\prime}$ $\iff$ $q^{-1}(B)\cap\mathcal{A}^{\prime}=(q^{-1}(B)\cap\mathcal{A}^{\prime\prime})\cap\mathcal{A}^{\prime}\in\tau^{\prime}$ $\stackrel{{\scriptstyle\textrm{construction of $\tau^{\prime\prime}$}}}{{\Longrightarrow}}$ $q^{-1}(B)\cap\mathcal{A}^{\prime\prime}\in\tau^{\prime\prime}$ $\iff$ $B\in\tau_{q}^{\prime\prime}$). This completes the proof of the claim.
@@ 6. Conclusion and questions
For a space $X$, we have seen (Theorem 3.3) that certain hyperspaces $\mathcal{J}\subset Cl(X)$ of $X$ can be described as quotients of function spaces $\mathcal{F}\subset X^{Y}$ in a natural way. Following this, we have discussed the concrete realization of certain preferred function space topologies (Theorem 4.2), the metrization of compact-subset hyperspaces (Theorems 4.6 and 4.9), and the existence of $\tau_{p}$-compatible function space topologies with $\tau_{v}$-compatible quotients (Theorem 5.1).
In addition to Questions 1 and 2, we have the following interesting questions.
Question 3.
Let $X,Y$ be spaces, $\mathcal{F}\subset X^{Y}$ a $q$-full subset, $Z\subset Y$, and $Cl_{Z}(X)\subset(Cl_{Y}(X),\tau)$ a $\tau$-closed subset, in which case a quotient map $q:(\mathcal{F},\widetilde{\tau})\rightarrow(Cl_{Y}(X),\tau)$ restricts to a quotient map $q:(\mathcal{F}_{Z},\widetilde{\tau})\subset X^{Y}\rightarrow(Cl_{Z}(X),\tau)$, where $\mathcal{F}_{Z}:=q^{-1}(Cl_{Z}(X))\cap\mathcal{F}$. If $X$ is a metric space, can we find a Lipschitz retraction
@@
Question 4.
Let $X$ be a space and $Y$ a set. From Definition 3, for which $Y$ (of smallest possible cardinality) does the equality $Cl_{Y}(X)=Cl(X)$ hold? Also, what is the cardinally-smallest $q$-full subset $\mathcal{F}\subset X^{Y}$?
Question 5 (Path representation problem).
Let $X$ be a space, $Y$ a set, $\mathcal{F}\subset X^{Y}$ a $q$-full subset, and $\tau_{\pi}$ a swrc-topology on $\mathcal{F}$. It is clear that a path $\eta:[0,1]\rightarrow(\mathcal{F},\tau_{\pi})$ gives a path $q\circ\eta:[0,1]\stackrel{{\scriptstyle\eta}}{{\longrightarrow}}\mathcal{F}\stackrel{{\scriptstyle q}}{{\longrightarrow}}(Cl_{Y}(X),\tau_{\pi q})$, since the composition of continuous maps is continuous. Conversely, (when) does every path $\gamma:[0,1]\rightarrow(Cl_{Y}(X),\tau_{\pi q})$ come from (or lift as $\gamma=q\circ\eta$ to) a path $\eta:[0,1]\rightarrow(\mathcal{F},\tau_{\pi})$?
Whenever the answer to this question is positive, every path $\gamma:[0,1]\rightarrow(Cl_{Y}(X),\tau_{\pi q})$ is expressible in the form
$\gamma(t)$ $=q\circ\eta(t)=q\big(\eta(t)(Y)\big)=cl_{X}\eta(t)(Y)=cl_{X}\{\eta(t)(y):y\in Y\}$
$\equiv cl_{X}{\{\gamma_{r}(t):r\in\Gamma\}},~~\textrm{~}\forall t\in[0,1],$
for a set of paths $\{\gamma_{r}:[0,1]\rightarrow X\}_{r\in\Gamma}$ in $X$. Of course, by the axiom of choice, any path $\gamma:[0,1]\rightarrow(Cl_{Y}(X),\tau_{\pi q})$ can be written as
@@
By the above paragraph, Question 5 is relevant to the “path representation” problem considered in [akofor2024] and therefore relevant to [akofor2024, Question 5.1], even though the current topology $\tau_{\pi q}$ on $BCl_{Y}(X)\subset Cl_{Y}(X)$ is in general not the same as the $d_{H}$-topology. Since $\gamma^{-1}(\mathcal{O})=\eta^{-1}(q^{-1}(\mathcal{O}))$ for any open set $\mathcal{O}\subset(Cl_{Y}(X),\tau_{\pi q})$, if $\tau_{\pi}=q^{-1}(\tau_{\pi q})$, then $\gamma$ is continuous iff $\eta$ is continuous (but in general, if we have proper containment $q^{-1}(\tau_{\pi q})\subsetneq\tau_{\pi}$, then $\gamma=q\circ\eta$ can be continuous even when $\eta$ is not continuous). Therefore, for $\gamma=q\circ\eta$ to be continuous, we only need a $q^{-1}(\tau_{\pi q})$-continuous (not necessarily a $\tau_{\pi}$-continuous) selection
@@
In particular, if $Y$ is finite and $X^{Y}$ is Hausdorff, then by [Vesely1991, Theorem 3.3], the continuous selection $\eta$ always exists, in which case, every path $\gamma:[0,1]\rightarrow Cl_{Y}(X)=FS_{|Y|}(X)$ has the form $\gamma(t)=\{\gamma_{r}(t):r\in\Gamma\}$, for paths $\gamma_{r}:[0,1]\rightarrow X$.
Acknowledgments
This manuscript has been improved using feedback from referees and editors of The Houston Journal of Mathematics.